EXERCISES the indicated operations, keeping the appropriate number of significant figures: 43.897+26.3 (4365) x (79) b) 230.04-2.3x10 d) 23.4x10 /8.459x10 a) For the following data sets find the mean x, and standard deviation of the me sure to report your results with the correct number of significant figures. 2) ean ơm, making a) b) 4.594; 4642; 4.613; 4.623 39.95; 40.10; 40.20; 40.05; 40.00; 39.98 (all given in meters) (all given in hours) 3) Find an expression for sa in terms of a, b, c, öb, and oc for the following cases: a) b+2c Given the information below and the accompanying formulas, determine the uncertainty associated with the derived quantity 4) m (34.5+0.2)kg (2.45 0.03) m/s find ptop a) p mv; b) K (12)2 The following table represents data for distance vs. time. a) b) Determine the slope, intercept and tncir associated uncertainties. 5) Using the guidelines described above, generate a graph of the data with titles, labels, error bars etc. distance (m) 3.5 ± 0.5 5.6 t 0.5 7.9 t 0.8 9.6+ 0.8 11.9+ 0.9 13.6+ 1.1 15.8 1.0 Time(s)
Answer #1

Solution 1)

A) 43.897+ 26.3 = 70.19= 70.2

B) 230.04-2.3*10=207.04

C) 4365*10=43650= 4.3 *10^4

D) 23.4*10^-2/8.459*10^-1 = 0.2766= 0.28

======

Acc to HOMEWORKLIB RULES, I can answer only one at a time. Good luck..pls rate